INFRARED DIODE LASER SPECTROSCOPY OF THE P=O (ν_{2}) FUNDAMENTAL BAND OF CIS-HOPO

\underline{PAUL J. O’SULLIVAN}, PETER A. HAMILTON and PAUL B. DAVIES, Department of Chemistry, University of Cambridge, Lensfield Road, Cambridge CB2 1EW, U.K..

The P=O (ν_{2}) stretching fundamental of the transient molecule cis-HOPO has been measured using high resolution infrared diode laser absorption spectroscopy. The molecule was formed in a flow system by reacting white phosphorus vapour and the products of a microwave discharge through a mixture of hydrogen and oxygen. Over 300 lines have been assigned between 1242cm^{−1} and 1267cm^{−1}. The band has the structure of an a-type transition of an asymmetric rotor. Most of the assigned lines were included in a least squares fit using Watson’s A-reduced Hamiltonian. The band origin is 1258.540021(12)cm^{−1}. The rotational constants are in satisfactory agreement with those from Density Functional Theory and with those predicted from a comparison with the nitrogen analogue, cis-HONO:

\begin{matrix}A_{0} = & 1.255036(55), & B_{0} = & 0.3188211(99), & C_{0} = & 0.2533085(91) \\ A_{1} = & 1.250424(55), & B_{1} = & 0.3181419(96), & C_{1} = & 0.2526089(96) \end{matrix}

Many of the remaining, unfitted lines are perturbed, particularly around K_{a}=7. The source of the perturbation will be considered.